A manufacturer of window frames knows from long experience that 5 percent of the production will have some type of minor defect that will require and adjustment. What is the probability that in a sample of 20 window frames:
a. None will need adjustment?
b. At least one will need adjustment?
c. More than two will need adjustment?
I know this is binomial but I am still having trouble solving the problem. Could some please help me and Thank you very much.

Answer by stanbon(75887)

A manufacturer of window frames knows from long experience that 5 percent of the production will have some type of minor defect that will require and adjustment. What is the probability that in a sample of 20 window frames:
a. None will need adjustment?
P(x = 0) = 20C0(0.05)^0*0.95^20 = 0.95^20 = 0.3585
--------------------
b. At least one will need adjustment?
P(1<= x <=20) = 1 - P(x = 0) = 1-3585 = 0.6415
--------------------
c. More than two will need adjustment?
P(3<= x <=20) = 1 - P(0<=x <=2) = 0.0755
=========
Cheers,
Stan H.
=========